Roberts demand curve for good X is given by the equation X 100 - 2Px 5 points) a. What is the elasticity of demand at the point X-20, Px -40 (5 points) b. If price falls from Px-40 to Py -35, what happens to total spending for X and what does this imply about the elasticity of demand? 5 points) c. Compute the elasticity to verify the answer.
Answer #1

A.For calculating elasticity,we will first differentiate with respect to price i.e, dX/dPx.

dX/dPx=(-2)

Now,for elasticity we know the formula is given as-

(dX/dPx).(Px/X)

=(-2).(40/20)

=(-4).

B. The elasticity declines withwthe change in prices.

C. In order to compute elasticity at Px=35,we'll first derive the quantity of X at this price i.e, X=100-2(35)=100-70=30.

Now elasticity is given as-

(dX/dPx).(Px/X)

=(-2).(35/30)=(-3.5).

Thus,the elasticity decline with the change in prices.